what is the sum of thrice a number and two equals 6 less than four times the number. find the number.

Answer by mananth(16946)

let number be x
3x+2=4x-6
-2
3x+2-2=4x-2-6
3x=4x-8
-4x
3x-4x=-8
-x=-8
multiply by -1
x=8 the number
...
m.ananth@hotmail.ca